# GENESIS: A Generative Model of Episodic–Semantic Interaction

### v1

@@ 3 Discussion
The GENESIS model makes several novel predictions that can be tested empirically. First, because episodic memories are encoded through the Cortical-VAE, the quality of recall should depend systematically on the representational capacity of semantic regions during encoding. Neural or behavioral markers of reduced cortical capacity (e.g., divided attention, low salience, strong expectation) should yield less detailed but more prototypical recollections—mirroring the compression–fidelity trade-off described by rate–distortion theory (Shannon et al.,, 1959; D’Amato et al.,, 2025). Second, the model predicts distinct memory failures depending on whether compression occurs in the Cortical-VAE or Hippocampal-VAE. Reduced cortical capacity should produce gist-based distortions, reflecting stronger semantic influence, whereas reduced hippocampal capacity should impair retrieval success without altering the perceptual quality of reconstructions. These effects could be tested by manipulating representational load or inhibition separately in cortical and hippocampal regions. Finally, the model predicts that replay and generative recombination can give rise to novel episodic constructions—such as imagined future events—whose plausibility and novelty are shaped by the statistical structure of the learned semantic space. Accordingly, neural or behavioral measures of creative recombination (e.g., in episodic future thinking) should reveal systematic effects of semantic similarity and compression on the balance between familiarity and originality in generated episodes.
This study has several limitations that open promising avenues for future research. First, for simplicity, the Cortical-VAE was pre-trained rather than co-trained with the episodic system. In line with the complementary learning systems (CLS) framework, episodic retrieval and replay from the Hippocampal-VAE could contribute to the gradual semantization of episodic traces in the Cortical-VAE, potentially implemented through generative replay (Stoianov et al.,, 2022; Shin et al.,, 2017) engaging both systems in a bidirectional training loop.
@@
Third, episodic memory was modeled as a hybrid system combining a generative Hippocampal-VAE with a passive retrieval-augmented memory (RAG). While this ensured conceptual clarity, real episodic memories are dynamic—subject to consolidation, reconsolidation, and decay. Future extensions could make the RAG component adaptive through continual replay and re-encoding, emulating updating, consolidation, and forgetting. The decay of episodic memories over time could produce gist-based distortions during reconstruction, over and above those induced by limited-capacity encoding discussed in this study.
@@
Fourth, future implementations could adopt more biologically realistic methods to form episodic memories, such as those based on behavioral time scale synaptic plasticity (Li and Roxin,, 2023; Wu and Maass,, 2025) or modern Hopfield networks (Ramsauer et al.,, 2020; Spens and Burgess,, 2024) and study how they organize experiences into cognitive maps (Behrens et al.,, 2018; George et al.,, 2021; Bottini and Doeller,, 2020; Schapiro et al.,, 2017; Van de Maele et al.,, 2024).
@@
Fifth, our simulations followed classical paradigms in which the model is externally instructed what to memorize (e.g., all items in a list). In natural settings, however, organisms autonomously decide when and what to store based on factors such as event segmentation, surprise, and saliency (Zacks,, 2020; Kumar et al.,, 2023). Future work could endow GENESIS with mechanisms for autonomous event boundary detection and memory updating (Lu et al.,, 2022).
@@
Sixth, recall dynamics were simplified. In the serial recall task, we did not implement a stopping rule, although recall empirically tends to terminate when confidence declines (Atkinson and Shiffrin,, 1968; Kahana,, 2020). Future implementations could include a confidence-based termination mechanism, where recall ceases when query–key similarity drops below threshold.
@@
Seventh, we did not address primacy effects, which may arise from stronger encoding or context reinstatement for early items (Howard and Kahana,, 2002). GENESIS could model these by allowing encoding capacity to vary with attention and novelty.
@@
Eighth, GENESIS currently abstracts away from the temporal structure of replay. Empirical data show that episodic retrieval involves temporally compressed reactivation (Wimmer et al.,, 2020; Howard,, 2018), with latency scaling with temporal distance (Hacker,, 1980; Muter,, 1979). Future studies could map cosine dissimilarities between temporal embeddings onto retrieval latency to model temporal dynamics of mental search and mental time travel (Suddendorf and Corballis,, 2007; Hills et al.,, 2015). Finally, future versions of GENESIS could test whether deficits in semantic (e.g., semantic dementia) or episodic (e.g., Alzheimer’s disease) processing can be explained in terms of capacity limitations (Patterson et al.,, 2007).
@@ 4.1 Cortical-VAE architecture and training
The Cortical-VAE was implemented as a conditional $\beta$-VAE (Higgins et al.,, 2017; Burgess et al.,, 2018; Sohn et al.,, 2015), trained to reconstruct colored MNIST digits ($x\in\mathbb{R}^{3\times H\times W}$). The dataset contained 50,000 training and 10,000 test samples, with each digit class evenly distributed across five colors. Images were resized to $H=W=32$. The encoder consisted of $m=3$ convolutional downsampling blocks, each with stride 2, kernel size 3, a residual block, and batch normalization, progressively reducing the input to spatial dimensions $h=H/2^{m}=4,\ w=W/2^{m}=4$. The decoder mirrored this structure with symmetric upsampling blocks. The latent representation was a Gaussian feature map $z\in\mathbb{R}^{16\times 4\times 4}$. Conditioning was implemented through two learnable embeddings, a digit embedding $e_{S}\in\mathbb{R}^{64}$ and a color embedding $e_{C}\in\mathbb{R}^{64}$, which were injected into the decoder via FiLM layers (Perez et al.,, 2018) together with the latent feature map $z$. The complete architecture had approximately 561k parameters. Training was performed with AdamW (learning rate 0.0025), batch size 64, and 10 epochs. Binary cross-entropy was used as the reconstruction loss. To impose a limited-capacity bottleneck, we applied the $\beta$-VAE objective with $\beta=100$ and a warm-up schedule of 3,000 iterations to gradually reach the target capacity $C$. The loss function was
@@ 4.5 Quantifying diversity in gist-based distorted image samples
Diversity was assessed using the Vendi score (VS; (Friedman and Dieng,, 2022)), which measures the self-similarity of a set of embeddings $v=[v_{1},v_{2},\ldots,v_{N}]$:

### v2

@@ 3 Discussion
The GENESIS model makes several novel predictions at the cognitive level that can be tested empirically. First, because episodic memories are encoded through the Cortical-VAE, the quality of recall should depend systematically on the representational capacity of semantic regions during encoding. Neural or behavioral markers of reduced cortical capacity (e.g., divided attention, low salience, strong expectation) should yield less detailed but more prototypical recollections—mirroring the compression–fidelity trade-off described by rate–distortion theory (Shannon and others, 1959; D’Amato et al., 2025). Second, the model predicts that the hippocampus forms “indexes,” in line with indexing theory (Teyler and DiScenna, 1986), by compressing latent cortical embeddings through the Hippocampal-VAE and integrating them with additional embeddings, such as temporal representations and potentially others (e.g., spatial context embeddings). Future studies could examine whether the hippocampus implements this form of compression or whether indexes are constructed through alternative mechanisms. Third, the model predicts distinct memory failures depending on whether compression occurs in the Cortical-VAE or Hippocampal-VAE. Reduced cortical capacity should produce gist-based distortions, reflecting stronger semantic influence, whereas reduced hippocampal capacity should impair retrieval success without altering the perceptual quality of reconstructions. These effects could be tested by manipulating representational load or inhibition separately in cortical and hippocampal regions. Another prediction is that certain tasks, such as the “old/new” recognition task, require synergistic contributions of cortical processes (to encode the item) and hippocampal processes (to perform the query–key match). Crucially, however, successful performance does not necessarily require retrieval of the associated value. This prediction could be tested by designing experiments that dissociate the cognitive and neural signatures of query–key matching from those of full value retrieval. A further prediction is that in certain cases, the task could be shortcut by the Cortical-VAE computing a coarse-grained sense of “familiarity” based on the similarity of the item to its learned semantic representations, without full episodic retrieval—but this sense of familiarity would lack episodic details. Finally, the model predicts that replay and generative recombination can give rise to novel episodic constructions—such as imagined future events—whose plausibility and novelty are shaped by the statistical structure of the learned semantic space. Accordingly, neural or behavioral measures of creative recombination (e.g., in episodic future thinking) should reveal systematic effects of semantic similarity and compression on the balance between familiarity and originality in generated episodes.
This study has several limitations that open promising avenues for future research. First, for simplicity, the Cortical-VAE was pre-trained rather than co-trained with the episodic system. However, this account does not capture the richer interplay between memory systems during consolidation, in which the hippocampus guides the reorganization of semantic information in the cortex (Squire, 1992; Squire et al., 2015; Nadel and Moscovitch, 1997; Moscovitch et al., 2005). The complementary learning systems framework proposes that the fast-learning hippocampal system replays experiences to train the slower-learning cortical system, leading to the gradual semanticization of episodic traces (McClelland et al., 1995). Relatedly, the transformation hypothesis suggests that hippocampally dependent episodic memories evolve into semantic or gist-like representations supported by cortical structures; although the hippocampus becomes unnecessary for retrieving semantic memories, it remains critical for episodic recollection (Winocur and Moscovitch, 2011). While GENESIS does not currently implement these mechanisms, they could be incorporated by using episodic retrieval and replay from the Hippocampal-VAE, as illustrated in Section 2.4, to train the Cortical-VAE via generative replay (Stoianov et al., 2022; Shin et al., 2017). Such an extension would establish a bidirectional interaction between episodic and semantic systems, accounting for both semanticization and the continued reliance on the Hippocampal-VAE for episodic retrieval even after semantic knowledge has formed.
@@
Third, episodic memory was modeled as a hybrid system combining a generative Hippocampal-VAE with a passive retrieval-augmented memory (RAG). While this ensured conceptual clarity, real episodic memories are dynamic—subject to consolidation, reconsolidation, and decay (Hupbach et al., 2007; Roediger and Butler, 2011; Gershman et al., 2013; Zaki et al., 2025). Future extensions could make the RAG component adaptive through continual replay and re-encoding, emulating updating, consolidation, and forgetting. The decay of episodic memories over time could produce gist-based distortions during reconstruction, over and above those induced by limited-capacity encoding discussed in this study (Schacter et al., 2011).
@@
Fourth, future implementations could adopt more biologically realistic methods to form episodic memories, such as those based on behavioral time scale synaptic plasticity (Li and Roxin, 2023; Wu and Maass, 2025) or modern Hopfield networks (Ramsauer et al., 2020; Spens and Burgess, 2024) and study how they organize experiences into cognitive maps (Behrens et al., 2018; George et al., 2021; Bottini and Doeller, 2020; Schapiro et al., 2017; Van de Maele et al., 2024). Relatedly, our current implementation relies on machine learning tools such as VAEs, RAG, and query–key–value mechanisms. These capture key computational principles we consider central—compression, flexible retrieval, and episodic indexing—but only in an abstract form that departs from biological implementation. For example, key–value memory architectures, which distinguish stored content (values) from retrieval cues (keys), are increasingly recognized as important, yet future work should explore biologically grounded realizations of these mechanisms (Kozachkov et al., 2023; Tyulmankov et al., 2021; Gershman et al., 2025).
@@
Fifth, our simulations followed classical paradigms in which the model is externally instructed what to memorize (e.g., all items in a list). In natural settings, however, organisms autonomously decide when and what to store based on factors such as event segmentation, surprise, and saliency (Zacks, 2020; Kumar et al., 2023). Future work could endow GENESIS with mechanisms for autonomous event boundary detection and memory updating (Lu et al., 2022). Relatedly, future extensions could enable GENESIS to store not only episodes but also their associated utility, potentially allowing the model to capture episodic influences on decision-making (Nicholas et al., 2025; Nicholas and Mattar, 2026; Zhou et al., 2024; Bornstein et al., 2017; Mason et al., 2025).
@@
Sixth, recall dynamics were simplified. In the serial recall task, we did not implement a stopping rule, although recall empirically tends to terminate when confidence declines (Atkinson and Shiffrin, 1968; Kahana, 2020). Future implementations could include a confidence-based termination mechanism, where recall ceases when query–key similarity drops below threshold.
@@
Seventh, we did not address primacy effects, which may arise from stronger encoding or context reinstatement for early items (Howard and Kahana, 2002). GENESIS could model these by allowing encoding capacity to vary with attention and novelty.
@@
Eighth, GENESIS currently abstracts away from the temporal structure of replay. Empirical data show that episodic retrieval involves temporally compressed reactivation (Wimmer et al., 2020; Howard, 2018), with latency scaling with temporal distance (Hacker, 1980; Muter, 1979). Future studies could map cosine dissimilarities between temporal embeddings onto retrieval latency to model temporal dynamics of mental search and mental time travel (Suddendorf and Corballis, 2007; Hills et al., 2015). It would also be possible to model sequential replay in GENESIS as a cascade of episodic retrievals, by chaining multiple query–key–value matches to generate context-dependent, temporally ordered reactivations of past events (Zhou et al., 2026; Stoianov et al., 2022).
Ninth, future developments of GENESIS should pursue a more detailed mapping between its computational components and the neurobiological mechanisms underlying semantic and episodic memory. For simplicity, we labeled the components of GENESIS according to the classical distinction between a fast-learning hippocampal episodic system and a slow-learning cortical semantic system. However, this dichotomy is overly simplistic. Converging evidence indicates that semantic memory also depends on medial temporal lobe regions (Binder and Desai, 2011; Renoult et al., 2019), while episodic memory critically engages neocortical regions (Addis, 2018; Rugg and Vilberg, 2013; Schacter et al., 2007). Accordingly, the present labels should be understood as heuristic rather than anatomical commitments. A key direction for future work is to clarify how the proposed components of GENESIS map onto interactive hippocampal–cortical dynamics supporting semantic and episodic memory.
Relatedly, semantic and episodic memory formation likely involve richer dynamics than currently captured by GENESIS. Contrary to the traditional view of a fast-learning hippocampal system and a slower-learning cortical system, evidence suggests that memory engrams can form rapidly in posterior parietal cortex, in a manner that appears partly hippocampus-independent (Brodt et al., 2016, 2018). Moreover, recent theoretical accounts propose that multiple representations of an episodic memory may be encoded simultaneously at different levels of granularity—from precise, context-specific details to more generalized, gist-like representations—and that semantic knowledge about specific events may form concurrently with the corresponding episodic memory rather than only afterward (Gilboa and Moscovitch, 2021; Tarder-Stoll et al., 2026). These complexities fall beyond the current scope of GENESIS but represent important directions for future research.
Finally, future versions of GENESIS could test whether deficits in semantic (e.g., semantic dementia) or episodic (e.g., Alzheimer’s disease) processing can be explained in terms of capacity limitations (Patterson et al., 2007).
@@ 4.1 Cortical-VAE architecture and training
The Cortical-VAE was implemented as a conditional $\beta$-VAE (Higgins et al., 2017; Burgess et al., 2018; Sohn et al., 2015), trained to reconstruct colored MNIST digits ($x\in\mathbb{R}^{3\times H\times W}$). The dataset contained 50,000 training and 10,000 test samples, with each digit class evenly distributed across five colors. Images were resized to $H=W=32$. The encoder consisted of $m=3$ convolutional downsampling blocks, each with stride 2, kernel size 3, a residual block, and batch normalization, progressively reducing the input to spatial dimensions $h=H/2^{m}=4,\ w=W/2^{m}=4$. The decoder mirrored this structure with symmetric upsampling blocks. The latent representation was a Gaussian feature map $z\in\mathbb{R}^{16\times 4\times 4}$. Conditioning was implemented through two learnable embeddings, a digit embedding $e_{S}\in\mathbb{R}^{64}$ and a color embedding $e_{C}\in\mathbb{R}^{64}$, which were injected into the decoder via FiLM layers (Perez et al., 2018) together with the latent feature map $z$. The complete architecture had approximately 561k parameters. Training was performed with AdamW (learning rate 0.0025), batch size 64, and 10 epochs. Binary cross-entropy was used as the reconstruction loss. To impose a limited-capacity bottleneck, we applied the $\beta$-VAE objective with $\beta=100$ and a warm-up schedule of 3,000 iterations to gradually reach the target capacity $C$. The loss function was
@@ 4.5 Quantifying diversity in gist-based distorted image samples
Diversity was assessed using the Vendi score (VS; (Friedman and Dieng, 2022)), which measures the self-similarity of a set of embeddings $v=[v_{1},v_{2},\ldots,v_{N}]$: